OpenStudy (lilcathyx):

all i have is a question because i forgot how to do it. its been a while since i done proportions and i was wondering how do you know when 2 objects are proportional ? for example a table is the length of 30 in and there's another table that is length of 15 in. prove that these two objects are proportional.

OpenStudy (aravindg):

You see by proportional it means they have the same ratio. 1/2=2/4=4/8

OpenStudy (aravindg):

So if someone says ratio of x:y is 1/2 and ratio of w:v is 2/4 Both are proportional to each other.

OpenStudy (lilcathyx):

ohh and how do i prove that 2 objects are similar ? sorry i haven't done these in a while and i forgot

OpenStudy (aravindg):

Usually we have figures like rectangle. Let length bread ratio be 1/2. If we construct another rectangle with ratio 2:4. Then both these are proportional. Geometrically this mean first figure has been maginified 2 times.

OpenStudy (lilcathyx):

ohh okk. but umm ok look i did a model of my table and i have to " Write a proof demonstrating how your model is similar to the object you chose."

OpenStudy (lilcathyx):

and im stuck there.

OpenStudy (aravindg):

The question is not so clear. Can you elaborate?

OpenStudy (lilcathyx):

i did a model of my table. my table is round. the round part of my table measures 30 inches and in my model i did it by half (which means my model round part measures 15 inches). now i have to prove how my model is similar to my table.

OpenStudy (lilcathyx):

and i dont know how to prove it. and then i got the idea of proportions but i dont think that will work..

OpenStudy (aravindg):

15*2=30 Essentially the bigger round table is maginifed 2 times than the smaller one.

OpenStudy (lilcathyx):

thats it ? thats how i know its similar? thats how i prove its similar ?

OpenStudy (aravindg):

Yeah. you could also do something like that for the legs. Maybe their height is also 2 times?

OpenStudy (lilcathyx):

yeah everything is 2 times bigger then. i did everything by half

OpenStudy (lilcathyx):

Thank youu :)

OpenStudy (aravindg):

yw :)